OpenStudy (anonymous):

I am probably asking the simpliest question ever but I suck at algebra and can't figure out what i am doing wrong. 124.8-9.4x=4.8x+32.5 and i have to find x

OpenStudy (anonymous):

think the answer is 6.5

OpenStudy (anonymous):

ok, but how........

OpenStudy (anonymous):

clear the annoying decimals and write \[1248-94x=48x+325\] so make your life easier

OpenStudy (anonymous):

then \[1248-325=48x+94x\] \[923=142x\]

OpenStudy (anonymous):

so \[x=\frac{923}{142}=6.5\]

OpenStudy (anonymous):

ok it just seems like i am missing a part and can't figure out what i am doing wrong. can you help me if i do one by myself and then let you know if i need help?

OpenStudy (anonymous):

steps clear?

OpenStudy (anonymous):

sure. i only cleared the decimals (multiplied both sides by 10) to make the calculation easier. you can work with the decimals if you like, but it is more annoying

OpenStudy (anonymous):

ok just give me a couple of minutes

OpenStudy (anonymous):

ok i don't think i am getting this. next probl is.62x-.65=9.75-2.63

OpenStudy (anonymous):

i get 3.382 w/ more numbers and that isn't the answer

OpenStudy (anonymous):

omg i got it!!!!

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

i am assuming that the problem you wrote is missing an x somewhere

OpenStudy (anonymous):

\[.62x-.65=9.75-2.63\]

OpenStudy (anonymous):

ugh but now i get the wrong answer for the next one. . k, i am going to show you what i am doing and you tell me what i am doing wrong ok?

OpenStudy (anonymous):

ok i will wait here

OpenStudy (anonymous):

you are so patient!!!

OpenStudy (anonymous):

5x+3=2x+18. So I moved the 3 over to the 18 and made it 5x=2x+15. then i moved the 2x over to the 5x which is then 3x=15 so the answer is 5 but the true answer is 3

OpenStudy (anonymous):

well you are right i think. lets check \[5x+3=2x+18\] \[5x-2x=18-3\] \[3x=15\] \[x=\frac{15}{3}=5\]

OpenStudy (anonymous):

so the "true answer" is not 3 and it is an easy check. if you replace x by 3 you get \[5\times 3 + 3=2\times 3 + 18\] \[15+3=6+18\] \[18=24\] which is wrong.

OpenStudy (anonymous):

but if you replace x by 5 you will get 28 on both sides so 5 is right and 3 is wrong

OpenStudy (anonymous):

misswrote numbers and the 18 should be a 12 now i have the right way to do it i think. thank you sosssoooo very very much!!!!!!!!!!!

OpenStudy (anonymous):

i think i get it now...i will continue working on my homework and will come back if i need more help

OpenStudy (anonymous):

ok. good luck

OpenStudy (anonymous):

ty